# Single-photon scattering in giant-atom topological-waveguide-QED systems

### v1

@@ II.3 Scattering spectra
By substituting $k_{\Delta}=-\pi/2$ (with $\Delta\approx 1.58J$) into Eqs. (22a)-(LABEL:approximated:phase), we can further obtain the Lamb shifts and decay rates of the giant atom for the AA- and AB-coupling configurations under different values of $\delta$ and $d$. The concrete expressions are summarized in Table II.3. It can be seen that the Lamb shifts and decay rates are periodically modulated by the coupling-point distance, with a period of $d=4$. When we take $d=4x+x_{0}$ (with $x_{0}=1$-$4$ for the AA-coupling case), the reflection coefficients in Eq. (23) become
@@
where the superscript represents the AA-coupling case. Equation (28) indicates that when $d=4x+1$, $34x+$ 3, and $44x+$ 4, the reflection spectra are characterized by the Lorentzian line shapes centered at $\Delta_{k}=-\Gamma_{e}$, $\Gamma_{e}$, and 00 with the linewidth of $2\Gamma_{e}$, $2\Gamma_{e}$, and $4\Gamma_{e}$, respectively. These results indicate that the approximation is valid. For $d=4x+2$, we have $\Delta_{L,k_{\Delta}}^{\text{AA}}=0$ and $\Gamma_{k_{\Delta}}^{\text{AA}}=0$. In this case, the giant atom is decoupled from the SSH waveguide, resulting in a complete transmission of the incident photon in the waveguide [the red solid curve and dashed curve with dots in Fig. 3(b)].
@@
For the AB-coupling configuration, both the Lamb shifts and decay rates in Table II.3 are nonzero for all $d=4x+x_{0}$ (with $x_{0}=0$-$3$) when $|\delta|=0.5$ and $k_{\Delta}=-\pi/2$. This makes the reflection spectra exhibit the Lorentzian line shapes centered at $\Delta_{k}=\Delta_{L,k_{\Delta}}^{\text{AB}}$ with spectrum linewidth of $\Gamma_{k_{\Delta}}^{\text{AB}}$, which is consistent with numerical results in Figs. 3(c) and 3(d). Specifically, when $d=4x+1$ and $\delta=0.5$ [the red solid curve in Fig. 3(c)], the Lamb shift and decay rate are given by $-\Gamma_{e}\cos\phi_{k_{\Delta},+}$ and $2\Gamma_{e}(1-\sin\phi_{k_{\Delta},+})$, respectively. When $d=4x+2$ and $\delta=-0.5$ [the red dashed curve with dots in Fig. 3(d)], the Lamb shift and decay rate are, respectively, given by $\Gamma_{e}\cos\phi_{k_{\Delta},+}$ and $2\Gamma_{e}(1-\sin\phi_{k_{\Delta},+})$. We find that the two reflection spectra have opposite Lamb shifts and the same decay rates. For $d=4x+3$ (with $\delta=0.5$) and $d=4x+4$ (with $\delta=-0.5$), the characteristic quantities of the reflection spectra exhibit opposite Lamb shifts and the same decay rates, as shown by the blue solid curve in Fig. 3(c) and the blue dashed curve with dots in Fig. 3(d).
To see the influence of the atom transition frequency and the dimerization parameter on the single-photon scattering, in Figs. 4(a), 4(c), and 4(e) we show the reflection coefficient as a function of $\Delta$ and $\Delta_{k}$ when $d$ takes a fixed value. For more details, we also plot in Figs. 4(b), 4(d), and 4(f) the profiles at some typical values of the detuning in the region of $\Delta\in(J,2J)$. We point out that the reflection spectra of AA-coupling configuration is independent of the sign of $\delta$. Thus, we only need to analyze the influence of $|\delta|$ on the scattering spectra. From Figs. 4(a) and 4(b) (AA-coupling case), the detuning $\Delta$ plays an important role in the single-photon scattering process. The linewidth $\Gamma_{k_{\Delta}}^{\text{AA}}$ of the reflection spectrum gradually narrows as $\Delta$ increases from $J$ to $1.58J$, and then progressively widens as $\Delta$ increases from $1.58J$ to $2J$. In particular, the reflection peak disappears completely when $\Delta\approx 1.58J$, as shown by the blue dashed curve in Fig. 4(b). This is because the giant atom is decoupled from the SSH waveguide. In addition, the scattering behavior is independent of the sign of $\delta$, which confirms our previous analyses.
Different from the AA-coupling case, the single-photon scattering behavior for the AB-coupling case is not only modulated by the detuning $\Delta$, but also depends on the sign of $\delta$. As shown in Figs. 4(c), 4(d), 4(e), and 4(f), the reflection spectra at $\delta=0.5$ and $\delta=-0.5$ indicate different dependencies on $\Delta$. In Figs. 4(c) and 4(d), when $\delta=0.5$, the linewidth $\Gamma_{k_{\Delta}}^{\text{AB}}$ of the reflection spectrum becomes wide as $\Delta$ increases from $J$ to $1.65J$, and then becomes narrow as $\Delta$ increases from $1.65J$ to $2J$. When $\Delta\approx 1.65J$, the linewidth of the reflection spectrum reaches its maximum value, as shown by the black dotted curve in Fig. 4(d). Meanwhile, the reflection peak is located at $\Delta_{k}=0$. For $\delta=-0.5$, as shown in Figs. 4(e) and 4(f), the dependence of the linewidths for the reflection spectra on the detuning $\Delta$ can be divided into three regions: (i) $\Gamma_{k_{\Delta}}^{\text{AB}}$ decreases monotonically with $\Delta$ in the region $\Delta\in(J,1.27J)$, (ii) $\Gamma_{k_{\Delta}}^{\text{AB}}$ increases monotonically with $\Delta$ in the region $\Delta\in[1.27J,1.77J)$, (iii) $\Gamma_{k_{\Delta}}^{\text{AB}}$ decreases monotonically with $\Delta$ in the region $\Delta\in[1.77J,2J)$. In particular, the reflection peak disappears completely when $\Delta\approx 1.27J$, as shown by the red solid curve in Fig. 4(f). This is also caused by the decoupling between the giant atom and the waveguide. In addition, the width of the reflection spectrum is the widest when $\Delta\approx 1.77J$, as shown by the purple dot-dashed curve in Fig. 4(f).
@@ III.1 Physical model and Hamiltonian
where $\Delta=\omega_{e}-\omega_{a}$ is the frequency detuning between the giant atoms and the cavity field. The atomic operators in Eq. (30) are defined by $\hat{\sigma}_{s=1,2}^{+}=\left|e\right\rangle_{ss}\left\langle g\right|$ and $\hat{\sigma}_{s}^{-}=\left|g\right\rangle_{ss}\left\langle e\right|$. Here, $\left|g\right\rangle_{s}$ and $\left|e\right\rangle_{s}$ are the ground state and excited state of the $s$th giant atom, respectively.
@@
$\hat{H}_{I,2}$ $=$ $g[(\mu_{1}\hat{a}_{n_{1}}^{{\dagger}}+\nu_{1}\hat{b}_{n_{1}}^{{\dagger}}+\mu_{2}\hat{a}_{m_{1}}^{{\dagger}}+\nu_{2}\hat{b}_{m_{1}}^{{\dagger}})\hat{\sigma}_{1}^{-}$ (31)
@@
Here, we assume that the $1(2)$th giant atom is coupled to the SSH waveguide at both the $n_{1(2)}$th and $m_{1(2)}$th unit cells. For simplicity, we consider that all the coupling points have the same coupling strength $g$. To discuss the photon scattering for different coupling configurations, in Eq. (31) we introduce the dimensionless parameters $\mu_{i}$ (for $i=1-4$) and $\nu_{i}$ to characterize the coupling between the two giant atoms and the SSH waveguide, where the coefficients $\mu_{i}$ and $\nu_{i}$ take either 00 or $1$. For example, for the ABAB-coupling configuration, we choose $\mu_{1(3)}=\nu_{2(4)}=1$ and $\mu_{2(4)}=\nu_{1(3)}=0$.
@@ III.2 Scattering probability amplitudes
where $c_{e1}$ $(c_{e2})$ is the probability amplitude corresponding to the single excitation stored in the 1(2)th giant atom, and $u_{k}(j)$ $[w_{k}(j)]$ has been introduced in Eq. (15). Similar to Sec. II, we can also obtain equations of motion for the probability amplitudes based on the stationary Schrödinger equation $\hat{H}_{\text{rot},2}\left|\Psi_{k}\right\rangle_{2}=\omega^{(\pm)}\left|\Psi_{k}\right\rangle_{2}$. In the case of $j<n_{1}$, $n_{1}<j<m_{1}$, $m_{1}<j<n_{2}$, $n_{2}<j<m_{2}$, and $j>m_{2}$, the probability amplitudes $u_{k}(j)$ and $w_{k}(j)$ corresponding to the upper and lower bands can be assumed to be
$u_{k}(j)$ $=$ $\left\{\begin{array}[]{cc}e^{ikj}+r_{2}e^{-ikj},&j<n_{1},\\ A_{1}e^{ikj}+B_{1}e^{-ikj},&n_{1}<j<m_{1},\\ A_{2}e^{ikj}+B_{2}e^{-ikj},&m_{1}<j<n_{2},\\ A_{3}e^{ikj}+B_{3}e^{-ikj},&n_{2}<j<m_{2},\\ t_{2}e^{ikj},&j>m_{2},\end{array}\right.$ (33f)
[…]
@@
$r_{2}^{(\pm)}\!\!$ =\!\!= $\frac{\Gamma_{1}^{(\pm)}\left(i\Delta_{k2}^{(\pm)}-i\Delta_{L2}^{(\pm)}-\Gamma_{2}^{(\pm)}/2\right)e^{2ikn_{1}+i\theta_{1}^{(\pm)}}\!\!+(1\leftrightarrow 2)+\Gamma_{1}^{(\pm)}\Gamma_{2}^{(\pm)}e^{2ikn_{2}+i\theta_{2}^{(\pm)}}}{2\left[\left(i\Delta_{k1}^{(\pm)}-i\Delta_{L1}^{(\pm)}-\Gamma_{1}^{(\pm)}/2\right)(1\leftrightarrow 2)-\left(\Gamma_{12}^{(\pm)}/2+ig_{12}^{(\pm)}\right)^{2}\right]},$
where $\Delta_{k1}^{(\pm)}=\Delta_{k2}^{(\pm)}=\Delta_{k}^{(\pm)}=\pm\omega_{k}-\Delta$ are the detunings between the upper/lower band propagating photons in the SSH waveguide and the giant atoms. Note that the notation ($1\leftrightarrow 2$) in Eqs. (35a) and (LABEL:r2) represent the term having the same form as the former term under the replacement of the subscripts $1\leftrightarrow 2$. We can check that the transmission and reflection amplitudes satisfy the relation $|t_{2}^{(\pm)}|^{2}+|r_{2}^{(\pm)}|^{2}=1$. It can be seen that the scattering amplitudes are determined by four characteristic quantities: the Lamb shifts $\Delta_{L1(2)}^{(\pm)}$, the individual decay rates $\Gamma_{1(2)}^{(\pm)}$ of the giant atom 1(2), the exchange interaction strengths $g_{12}^{(\pm)}$, and the collective decay rates $\Gamma_{12}^{(\pm)}$ of the two giant atoms, which are defined as
@@
$+\mu_{1}\nu_{4}\cos[k(d_{21}+d_{1}+d_{2})+\phi_{\mu}]$
$+\mu_{1}\nu_{4}\sin[k(d_{21}+d_{1}+d_{2})+\phi_{\mu}]$
In Eqs. (36a)-(36h), $d_{1(2)}=m_{1(2)}-n_{1(2)}$ is the coupling-point distance of the $1$$(2)$th giant atom and $d_{21}=n_{2}-m_{1}$ is the distance between the $2$nd coupling point of the giant atom 1 and the $1$st coupling point of the giant atom 2. To get more compact expressions, we have introduced $\phi_{\mu}=-\phi_{\nu}=-\phi_{k}$ in Eqs. (36e) and (36f).
Table 2: The relationships between the reflection coefficients for different coupling configurations. Here, the superscript $\text{C}_{1}\text{C}_{2}\text{C}_{3}\text{C}_{4}$ denotes the coupling configuration with $\text{C}_{i}\in\{\text{A},\text{B}\}$ (for $i=1-4$) and the subscript $\pm$ ()minus-or-plus($\mp$)( ∓ ) represent the sign of the dimerization parameter $\delta$.
@@
In Eqs. (35a)-(36f), we consider that the incident photon frequency is $\omega_{k}$ with $k\in(-\pi,0)$ and takes different values of $\mu_{i}$ and $\nu_{i}$ for $i=1-4$. We can obtain the concrete forms of the reflection coefficients for 16 coupling configurations. The relations of these reflection coefficients under different coupling distances and dimerization parameters are listed in Table 2. By changing the value of $d$ and the sign of $\delta$, these 16 coupling configurations can be reduced into six coupling cases: AAAA, AAAB, AABA, ABBA, and AABB couplings. To explain the physical mechanism behind this phenomenon, we introduce the symmetric-antisymmetric (S-A) lowering operators $\hat{\sigma}_{\text{S,A}}^{-}=(\hat{\sigma}_{1}^{-}\pm\hat{\sigma}_{2}^{-})/\sqrt{2}$ Bello2019 . In this case, the two two-level giant atoms in the single-excitation subspace can be understood as a V-type three-level atom coupled to the SSH waveguide through the symmetric- and antisymmetric-state channels. For different coupling configurations, the symmetric and antisymmetric states are coupled to the SSH waveguide with different coupling strengths. In the S-A state representation, we introduce the characteristic parameters as
@@
Here, $\Delta_{\text{S(A)}}$ is the effective detuning between the propagating photons in the waveguide and the symmetric (antisymmetric) state, $\Gamma_{\text{S(A)}}$ is the individual decay rate of the symmetric (antisymmetric) state, $g_{\text{SA}}$ and $\Gamma_{\text{SA}}$ are the exchange interaction strength and collective decay rate, respectively. According to Eqs. (37a)-(37e), the transmission and reflection amplitudes in Eqs. (35a) and (LABEL:r2) can be reexpressed by $\Delta_{\text{S(A)}}$, $\Delta_{\text{S(A)}L}$, $\Gamma_{\text{S(A)}}$, $g_{\text{SA}}$, and $\Gamma_{\text{SA}}$.
@@ III.3 Scattering spectra
Based on the relationship indicated by the reflection coefficients in Table 2, we only need to analyze the scattering spectra of these six coupling configurations. In the following, we investigate the influence of the coupling configurations, frequency detuning, and sign of the dimerization parameter on the single-photon scattering in the weak-coupling regime. Similar to the method in Sec. II, we can replace the wave vector $k$ in Eqs. (37a)-(37e) with $k_{\Delta}$. Note that the single-photon scattering is also periodically modulated by the coupling-point distance (with a period of 4). Therefore, we take fixed values of $d$ in the following discussions.

### v2

@@ II.3 Scattering spectra
Under the same parameter conditions, the approximated reflection spectra coincide with the original reflection spectra.
[…]
@@
where the superscript represents the AA-coupling case. Equation (28) indicates that, when $d=4x+1$, $34x+$ 3, and $44x+$ 4, the reflection spectra are characterized by the Lorentzian line shapes centered at $\Delta_{k}=-\Gamma_{e}$, $\Gamma_{e}$, and 00 with the linewidth of $2\Gamma_{e}$, $2\Gamma_{e}$, and $4\Gamma_{e}$, respectively. These results indicate that the approximation is valid. For $d=4x+2$, we have $\Delta_{L,k_{\Delta}}^{\text{AA}}=0$ and $\Gamma_{k_{\Delta}}^{\text{AA}}=0$. In this case, the giant atom is decoupled from the SSH waveguide, resulting in a complete transmission of the incident photon in the waveguide [the red solid curve and dashed curve with dots in Fig. 3(b)].
[…]
To see the influence of the atom transition frequency and the dimerization parameter on the single-photon scattering, in Figs. 4(a), 4(c), and 4(e) we show the reflection coefficient as a function of $\Delta$ and $\Delta_{k}$ when $d$ takes a fixed value. For more details, we also plot in Figs. 4(b), 4(d), and 4(f) the profiles at some typical values of the detuning in the region of $\Delta\in(J,2J)$. We point out that the reflection spectra of AA-coupling configuration is independent of the sign of $\delta$. Thus we only need to analyze the influence of $|\delta|$ on the scattering spectra. From Figs. 4(a) and 4(b) (AA-coupling case), the detuning $\Delta$ plays an important role in the single-photon scattering process. The linewidth $\Gamma_{k_{\Delta}}^{\text{AA}}$ of the reflection spectrum gradually narrows as $\Delta$ increases from $J$ to $1.58J$ and then progressively widens as $\Delta$ increases from $1.58J$ to $2J$. In particular, the reflection peak disappears completely when $\Delta\approx 1.58J$, as shown by the blue dashed curve in Fig. 4(b). This is because the giant atom is decoupled from the SSH waveguide. In addition, the scattering behavior is independent of the sign of $\delta$, which confirms our previous analyses.
@@
Different from the AA-coupling case, the single-photon scattering behavior for the AB-coupling case is not only modulated by the detuning $\Delta$, but also depends on the sign of $\delta$. As shown in Figs. 4(c), 4(d), 4(e), and 4(f), the reflection spectra at $\delta=0.5$ and $\delta=-0.5$ indicate different dependencies on $\Delta$. In Figs. 4(c) and 4(d), when $\delta=0.5$, the linewidth $\Gamma_{k_{\Delta}}^{\text{AB}}$ of the reflection spectrum becomes wide as $\Delta$ increases from $J$ to $1.65J$ and then becomes narrow as $\Delta$ increases from $1.65J$ to $2J$. When $\Delta\approx 1.65J$, the linewidth of the reflection spectrum reaches its maximum value, as shown by the black dotted curve in Fig. 4(d). Meanwhile, the reflection peak is located at $\Delta_{k}=0$. For $\delta=-0.5$, as shown in Figs. 4(e) and 4(f), the dependence of the linewidths for the reflection spectra on the detuning $\Delta$ can be divided into three regions: (i) $\Gamma_{k_{\Delta}}^{\text{AB}}$ decreases monotonically with $\Delta$ in the region $\Delta\in(J,1.27J)$, (ii) $\Gamma_{k_{\Delta}}^{\text{AB}}$ increases monotonically with $\Delta$ in the region $\Delta\in[1.27J,1.77J)$ and (iii) $\Gamma_{k_{\Delta}}^{\text{AB}}$ decreases monotonically with $\Delta$ in the region $\Delta\in[1.77J,2J)$. In particular, the reflection peak disappears completely when $\Delta\approx 1.27J$, as shown by the red solid curve in Fig. 4(f). This is also caused by the decoupling between the giant atom and the waveguide. In addition, the width of the reflection spectrum is the widest when $\Delta\approx 1.77J$, as shown by the purple dot-dashed curve in Fig. 4(f).
@@ III.1 Physical model and Hamiltonian
where $\Delta=\omega_{e}-\omega_{a}$ is the frequency detuning between the giant atoms and the cavity field. The atomic operators in Eq. (30) are defined by $\hat{\sigma}_{s=1,2}^{+}=\left|e\right\rangle_{ss}\left\langle g\right|$ and $\hat{\sigma}_{s}^{-}=\left|g\right\rangle_{ss}\left\langle e\right|$. Here, $\left|g\right\rangle_{s}$ and $\left|e\right\rangle_{s}$ are the ground state and excited state of the giant atom $s$, respectively.
@@
$\hat{H}_{I,2}$ $=$ $g[(\mu_{1}\hat{a}_{n_{1}}^{{\dagger}}+\nu_{1}\hat{b}_{n_{1}}^{{\dagger}}+\mu_{2}\hat{a}_{m_{1}}^{{\dagger}}+\nu_{2}\hat{b}_{m_{1}}^{{\dagger}})\hat{\sigma}_{1}^{-}$
@@
[…]
@@ III.2 Scattering probability amplitudes
where $c_{e1}$ $(c_{e2})$ is the probability amplitude corresponding to the single excitation stored in the giant atom 1(2), and $u_{k}(j)$ $[w_{k}(j)]$ has been introduced in Eq. (15). Similar to Sec. II, we can also obtain equations of motion for the probability amplitudes based on the stationary Schrödinger equation $\hat{H}_{\text{rot},2}\left|\Psi_{k}\right\rangle_{2}=\omega^{(\pm)}\left|\Psi_{k}\right\rangle_{2}$. In the case of $j<n_{1}$, $n_{1}<j<m_{1}$, $m_{1}<j<n_{2}$, $n_{2}<j<m_{2}$, and $j>m_{2}$, the probability amplitudes $u_{k}(j)$ and $w_{k}(j)$ corresponding to the upper and lower bands can be assumed to be
$u_{k}(j)$ =$=\begin{cases}e^{ikj}+r_{2}e^{-ikj},&j<n_{1},\\ A_{1}e^{ikj}+B_{1}e^{-ikj},&n_{1}<j<m_{1},\\ A_{2}e^{ikj}+B_{2}e^{-ikj},&m_{1}<j<n_{2},\\ A_{3}e^{ikj}+B_{3}e^{-ikj},&n_{2}<j<m_{2},\\ t_{2}e^{ikj},&j>m_{2},\end{cases}$ (33a)
$w_{k}(j)$ ={±$=\begin{cases}\pm e^{-i\phi_{k}}e^{ikj}\pm e^{i\phi_{k}}r_{2}e^{-ikj},&j<n_{1},\\ \pm e^{-i\phi_{k}}A_{1}e^{ikj}\pm e^{i\phi_{k}}B_{1}e^{-ikj},&n_{1}<j<m_{1},\\ \pm e^{-i\phi_{k}}A_{2}e^{ikj}\pm e^{i\phi_{k}}B_{2}e^{-ikj},&m_{1}<j<n_{2},\\ \pm e^{-i\phi_{k}}A_{3}e^{ikj}\pm e^{i\phi_{k}}B_{3}e^{-ikj},&n_{2}<j<m_{2},\\ \pm e^{-i\phi_{k}}t_{2}e^{ikj},&j>m_{2}.\end{cases}$ (33b)
@@
$r_{2}^{(\pm)}\!\!$ =\!\!= $\frac{\Gamma_{1}^{(\pm)}\left(i\Delta_{k2}^{(\pm)}-i\Delta_{L2}^{(\pm)}-\Gamma_{2}^{(\pm)}/2\right)e^{2ikn_{1}+i\theta_{1}^{(\pm)}}\!\!+(1\leftrightarrow 2)+\Gamma_{1}^{(\pm)}\Gamma_{2}^{(\pm)}e^{2ikn_{2}+i\theta_{2}^{(\pm)}}}{2\left[\left(i\Delta_{k1}^{(\pm)}-i\Delta_{L1}^{(\pm)}-\Gamma_{1}^{(\pm)}/2\right)(1\leftrightarrow 2)-\left(\Gamma_{12}^{(\pm)}/2+ig_{12}^{(\pm)}\right)^{2}\right]},$ (35b)
where $\Delta_{k1}^{(\pm)}=\Delta_{k2}^{(\pm)}=\Delta_{k}^{(\pm)}=\pm\omega_{k}-\Delta$ are the detunings between the upper and lower band propagating photons in the SSH waveguide and the giant atoms. Note that the notation ($1\leftrightarrow 2$) in Eqs. (35a) and (35b) represents the term having the same form as the former term under the replacement of the subscripts $1\leftrightarrow 2$. We can check that the transmission and reflection amplitudes satisfy the relation $|t_{2}^{(\pm)}|^{2}+|r_{2}^{(\pm)}|^{2}=1$. It can be seen that the scattering amplitudes are determined by four characteristic quantities: the Lamb shifts $\Delta_{L1(2)}^{(\pm)}$, the individual decay rates $\Gamma_{1(2)}^{(\pm)}$ of the giant atom 1(2), the exchange interaction strengths $g_{12}^{(\pm)}$, and the collective decay rates $\Gamma_{12}^{(\pm)}$ of the two giant atoms, which are defined as
@@
[…]
@@
$+\mu_{1}\nu_{4}\cos[k(d_{21}+d_{1}+d_{2})+\phi_{\mu}]$
$+\mu_{1}\nu_{4}\sin[k(d_{21}+d_{1}+d_{2})+\phi_{\mu}]$
[…]
@@
Here, $\Delta_{\text{S(A)}}$ is the effective detuning between the propagating photons in the waveguide and the symmetric (antisymmetric) state, $\Gamma_{\text{S(A)}}$ is the individual decay rate of the symmetric (antisymmetric) state, and $g_{\text{SA}}$ and $\Gamma_{\text{SA}}$ are the exchange interaction strength and collective decay rate, respectively. According to Eqs. (37a)–(37e), the transmission and reflection amplitudes in Eqs. (35a) and (35b) can be reexpressed by $\Delta_{\text{S(A)}}$, $\Delta_{\text{S(A)}L}$, $\Gamma_{\text{S(A)}}$, $g_{\text{SA}}$, and $\Gamma_{\text{SA}}$.
@@ III.3 Scattering spectra
Based on the relationship indicated by the reflection coefficients in Table 2, we only need to analyze the scattering spectra of these six coupling configurations. In the following, we investigate the influence of the coupling configurations, frequency detuning, and sign of the dimerization parameter on the single-photon scattering in the weak-coupling regime. Similar to the method in Sec. II, we can replace the wave vector $k$ in Eqs. (37a)–(37e) with $k_{\Delta}$. Note that the single-photon scattering is also periodically modulated by the coupling-point distance (with a period of 4). Therefore, we take fixed values of $d$ in the following discussions.
Figure 5: Reflection coefficient $R_{2}$ for the AAAA-coupling configuration as a function of $\Delta_{k}$ and $\Delta$. The dimerization parameter is (a) $|\delta|=0.5$. The profiles of panel (a) at several values of $\Delta$ are shown in (b). Other parameters used are $g=0.01J$ and $d=2$.